X^2+21x-72 = 0 how to solve for x and y?
Answer by rapaljer(4671)

Before we start this problem, you need to be able to FACTOR A TRINOMIAL. If ANYONE does NOT know how to factor a trinomial, PLEASE send me an Email at rapaljer@scc-fl.edu so I can send you some help on this MOST important topic!!

x^2 +21x-72=0

This is a quadratic equation that factors, so
(x+24)(x-3)=0

There are two solutions:
x+24=0 so x =-24
x-3=0 so x= 3

I don't see a y in the problem. You can't find y if you can't see it!!!

Dr. Robert J. Rapalje